OpenStudy (anonymous):

Solve the following system: y = x + 4 5x + y = 16 (2, -6) (-2, 6) (6, 2) (2, 6)

OpenStudy (anonymous):

You can substitute x + 4 for y in the 2nd equation: 5x + y = 16 5x + (x + 4) = 16 Can you simplify that? @blueSurrency

OpenStudy (anonymous):

We can add 5x + x, which gives us 6x. 6x + 4 = 16 Now can you do the rest?

OpenStudy (anonymous):

It would mean a lot if you could just tell me the answer and show me how you got it.

OpenStudy (anonymous):

I got it @iGreen. thank you

OpenStudy (anonymous):

@blueSurrency I'm not supposed to give answers..I'm supposed to guide you to it.

OpenStudy (anonymous):

What answer did you come up with?

OpenStudy (anonymous):

@blueSurrency